OpenStudy (anonymous):

the length of a regulation court in the NBA is 94 feet, how many meters is this? PLs show work i dont get this question

OpenStudy (dumbcow):

1 meter = 100 cm = 100/2.54 inches = 100/(12*2.54) feet

OpenStudy (dumbcow):

or going the other way 1 foot = 12 in = 12*2.54 cm = (12*2.54)/100 meters